OpenStudy (anonymous):

Solve: 5050= 2(w*17+50*w+850)

OpenStudy (anonymous):

5050=2(17w+50w+850) 5050=2(67w+850) 5050=134w+1700 5050-1700=134w 3350=134w w=3350/134=25

OpenStudy (anonymous):

25 is the answer

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

ur welcum

OpenStudy (anonymous):

hey where r u frm?